# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys using DFPT

### v3

@@ Abstract
The elastic tensor provides valuable insight into the mechanical behavior of a material with lattice strain, such as disordered binary alloys. Traditional stress-strain methods have made it possible to compute elastic constants for high-throughput databases of ordered structures, as well as individually tailored alloy compositions. However, this approach depends on predetermined or iteratively-chosen strain tensors. This poses a significant challenge for systematic, composition-dependent studies of disordered materials requiring structural models with randomized atomic positions and possessing low symmetry. Density functional perturbation theory (DFPT) based on density functional theory provides a compelling alternative to stress-strain methods: it allows for an unbiased determination of the elastic tensor, as well as access to local field data derived from the underlying general response function framework. Despite its intrinsic flexibility and efficiency, DFPT has seen limited application to the study of disordered systems. At the same time, there is a growing need for expanded quantum mechanical data to improve predictive modeling of complex disordered material properties. Here we present results for the rigid-ion and relaxed-ion elastic tensors computed using DFPT, for a comprehensive set of structural refractory body-centered cubic (BCC) binary alloys of molybdenum (Mo), niobium (Nb), tantalum (Ta), and tungsten (W). For the first time, we have mapped the quantum-driven heterogeneity in elastic constants and shear modulus, and associated relaxation fields at each disordered structure lattice site, by computing the force response internal strain tensor ($\Lambda$) and displacement response internal strain tensors ($\Gamma$). Derived properties—the bulk modulus ($B$), shear modulus ($G$), Young’s modulus ($E$), Poisson’s ratio ($\nu$), Pugh’s ratio ($B/G$), Cauchy pressure and elastic anisotropy—are reported as a function of composition for all refractory binaries. The DFPT-computed mechanical properties data for the refractory binary alloys at systematically-varied Mo, Nb, Ta, and W compositions are in excellent agreement with available experimental data.
@@ 1 Introduction
The elastic tensor [1] and its elastic constant components give a complete description, in terms of deformation, of material response under external stress within the elastic limit. The elastic tensor can therefore serve as a valuable quantitative probe of elastic strain effects associated with changes in chemical composition, and their impact on mechanical [2, 3, 4, 5, 6], electronic [7, 8], thermodynamic [9, 10], and structural [11, 12, 13] properties in these materials.
@@
Refractory alloys show great potential to meet the demands for high-temperature materials due to their high melting points and ability to maintain mechanical properties at temperatures higher than current Ni-based superalloys [14]. The refractory elements Nb, Ta, Mo, and W of groups V and VI comprise the BCC structural refractory metals family, characterized by high mechanical strength and melting temperatures [15]. Their structural refractory binary alloys MoNb, MoTa, MoW, NbTa, WNb and WTa form solid solutions of BCC crystal structure across their entire compositional space [16], and form an attractive starting point for designing more complex disordered alloys. However, elastic constant measurements are typically reported only for a limited number of alloy classes and compositions, and experiments often exhibit substantial disagreement even between values measured using the same technique and on the same material [17, 18]. Multiple factors can influence differences in reported experimental values, such as impurities and measurement temperature [19, 20, 21, 22]. Previous efforts have focused on the computational determination of elastic constants using density functional theory (DFT) for selected refractory binary alloys, using stress-strain methods [23, 24, 25, 26, 27]. While these investigations have provided valuable insights, they were limited to a few binary systems or computed for a narrow compositional range. The computational methodologies employed in these studies were primarily validated using experimental data for pure elements. Since the electronic interactions in alloy systems are inherently more complex than in their elemental counterparts, the direct application of methods validated for pure elements to alloy systems may introduce significant uncertainties, potentially compromising the accuracy and practical relevance of the predicted elastic properties for these alloys. These challenges are compounded when modeling disordered materials [28]. Here, we establish and validate a first-principles modeling framework based on DFT [29, 30] for performing computationally-efficient studies of elastic tensor and derived properties of disordered alloys, validated for systematically-varied Mo, Nb, Ta, and W binary alloy compositions. DFT, as implemented in state-of-the-art plane wave pseudopotential periodic electronic structure codes such as Abinit [31, 32], provides direct access to the computation of energies and stresses for obtaining the elastic tensors of the materials, with only atomic structure and nuclear charges as inputs.
@@
where $\bm{\sigma}$ is the $6\times 1$ stress tensor, the $C_{ij}$ are the elastic constants, and $\bm{s}$ is the $6\times 1$ strain tensor in Voigt notation [33].
In the traditional stress-strain approach, the elastic constants $C_{ij}$ are obtained by computing either the total energy or the stress of a crystal at a series of preselected strains [19, 34, 35, 36, 37]. The selection of strain sets can significantly influence the computed elastic constants, with variations exceeding 15% [38] in some cases. To mitigate this problem, high-throughput databases of ordered materials, such as the Materials Project [39], validate computed elastic constants by comparing results obtained from multiple strain sets [38]. In disordered alloys, such as the refractory binary alloys that are the focus of the present work, this approach is not practical due to the lack of crystal symmetry caused by intrinsic lattice strain.
An alternative and more general approach that avoids the need for tailored strains is based on analytical differentiation of the energy using linear response theory [40, 41, 42, 43]. In this method, the elastic tensor is obtained by expanding the total energy with respect to infinitesimal perturbation of the system using density functional perturbation theory (DFPT) [44].
@@
DFPT for the elastic tensor was initially implemented for norm-conserving pseudopotentials (NCPPs) [45] and later extended to the projector augmented-wave (PAW) method by Martin et al. [46]. DFPT has been successfully demonstrated using both NCPPs [47, 48, 49] and PAW pseudopotentials [50, 51, 52, 53, 54, 55]. Despite its intrinsic advantages, DFPT for elastic properties has been applied to only a limited number of structural alloys [56]. This work represents the first demonstration of DFPT for computing the elastic properties of the structural refractory disordered binary alloys of Mo, Nb, Ta and W.
Within the DFPT framework, the elastic tensor can be evaluated as the rigid-ion elastic tensor, $C_{ij}^{rigid}$, or as the relaxed-ion elastic tensor $C_{ij}^{relaxed}$ [40, 41, 42, 43]. $C_{ij}^{rigid}$ is computed with the relaxed initial structure, and strain as the perturbation; $C_{ij}^{relaxed}$ is a correction to the rigid-ion elastic tensor, with both strain and atomic displacements (nuclear relaxations) as the perturbations. Both $C_{ij}^{rigid}$ and $C_{ij}^{relaxed}$ provide essential information about the elastic response of a material. The relaxed-ion elastic tensor is the quantity that will correspond most closely to experiment.
Comparison of the rigid-ion and relaxed-ion results allows an assessment of the accuracy of the rigid-ion approximation, as well as enabling the mapping of the relaxation field of a disordered alloy. We define the nuclear relaxation elastic tensor $C_{ij}^{\text{nuclear}}$ as the difference between the corresponding rigid-ion and relaxed-ion quantities [43]:
@@
Non-zero nuclear-relaxation contributions in binary alloys are due to lattice distortions induced by the alloying elements, leading to forces between atoms in a finite strain state. Nuclear-relaxation contributions have previously been used to understand the effect of stress on piezoelectric coefficients and elastic constants by decomposing these properties into a purely electronic and nuclear-relaxation contributions resulting from the atomic rearrangements in the strained state [50, 51]. In the study by Zhang et al. [57], the nuclear-relaxation term of the piezoelectric coefficient was further decomposed into lattice point-specific relaxation contributions in order to elucidate the microscopic origin of the piezoelectric effect in $\alpha$-quartz. Despite the methodological similarities between the computation of relaxed-ion elastic constants and relaxed-ion piezoelectric properties, the specific contributions from nuclear relaxation to the elastic constants have not been previously reported. In this study, for the first time, microscopic insights into the influence of alloying on the mechanical behavior of refractory binary alloys are provided through the decomposition of the nuclear relaxation contributions. This approach enables the investigation of the spatial variation of elastic constants and derived elastic moduli within the crystal structure of disordered refractory binary alloys. This allows a deeper understanding of how local atomic environments affect the overall mechanical response of these complex systems. Hereafter, the terms “nuclear-relaxation contributions” and “relaxation values” will be used interchangeably.
@@
In the present work, we utilize DFT+DFPT to compute the single-crystal, relaxed-structure elastic tensors of the six BCC refractory binary alloys of (Mo, Nb, Ta, W)—MoNb, MoTa, MoW, NbTa, MoTa, and MoW—as a function of composition. The disordered alloys are modeled as supercells constructed using the special quasi-random structures (SQS) method [58]. The relaxed disordered structures are then used to compute the rigid-ion and relaxed-ion elastic tensors based on DFPT. We assess the effectiveness of DFPT in providing accurate elastic tensors through detailed comparison of elastic constants with available experimental data. The composition-dependent calculations serve as an important test of the ability of the SQS method to accurately model disorder in these alloys. Mechanical properties $B$, $G$, and $E$, the elastic anisotropy and $B/G$ and $C^{\prime}$ derived from the computed elastic moduli are reported as a function of composition for all refractory binaries. Computed relaxation fields for the Voigt-averaged elastic constants and shear modulus are presented for all binaries and discussed in terms of constituent element chemistry. The in-depth understanding of trends in elastic properties and anisotropy in the disordered refractory binary alloys gained in this work through the decomposition of elastic constants into electronic and nuclear contributions is expected to contribute to new microscopic alloy computational strategies for designing refractory alloys for high-temperature applications.
@@ 2 Computational Methods
We used the open-source, periodic plane wave pseudopotential electronic structure code Abinit [31, 32] for all DFT calculations, employing the Perdew, Burke, and Ernzerhof (PBE) generalized gradient approximation (GGA) exchange-correlation functional [59], with corresponding PBE norm-conserving pseudopotentials [60]. We note that while meta-GGA methods such as r2SCAN [61] and parameterized DFT+$U$ [62] approaches have seen growing popularity for computing the properties of ordered systems with 1–3 elements [63], the corresponding pseudopotentials are not currently available, and the accuracy of these functionals for complex disordered systems with 5+ elements has yet to be assessed. Since the present work is intended as a starting point for scaling up to larger refractory alloy systems with increased chemical complexity, the PBE functional is utilized throughout this work, to ensure a uniform level of theory across all alloys.
@@
The plane waves for each system were expanded up to a kinetic energy of 30 Ha, with each structure converged until the total residual differences in the potential were $<10^{-12}$ Ha, corresponding to convergence to $10^{-6}$ bohr for lattice constants, and atomic forces converged to <|±$<|\pm 0.01|$ mHa/bohr. Following careful convergence studies, we adopted a $\Gamma$-centered Monkhorst–Pack [64] special $k$-point grid with a density of 50 $k$-points per Å for all systems.
@@
Initial benchmark studies were performed on bulk Mo, Nb, Ta, and W in order to determine appropriate computational parameters for the binary alloy calculations. We computed the lattice constants, bulk moduli, cohesive energies, elastic constants, and (100), (110), (111) surface energies for elemental Mo, W, Nb, and Ta, and compared our results with available experimental data, using three different pseudopotentials: the FHI norm-conserving pseudopotentials ($d$ valence electrons only) tabulated at the Abinit website [65], pseudo-dojo [60] NCPPs ($s$, $p$ and $d$ valence electrons) available at [66], and projector augmented-wave (PAW) pseudopotentials [67] ($s$, $p$ and $d$ valence electrons), tabulated at the Abinit website [65]. Computed properties using the FHI NCPPs showed an average deviation of $\approx 1\%$ compared to available experiments, across all computed properties for Mo, Nb and Ta, but an average deviation of $\approx 9\%$ for W. For the pseudo-dojo NCPPs and PAW-based pseudopotentials, computed properties agreed well with experiment, with an average deviation of $\approx 1.5\%$ for all elements and across all properties. Detailed results are reported in [68]. Given the comparable performance of the pseudo-dojo NCPP and PAW pseudopotentials, the pseudo-dojo NCPPs were used throughout this work. Working within the pseudo-dojo framework [60] also provides a foundation for the custom construction of pseudopotentials based on emerging functionals, for computationally-consistent modeling of alloys of growing complexity.

### v4

@@ Abstract
The elastic tensor provides valuable insight into the mechanical behavior of a material with lattice strain, such as disordered binary alloys. Traditional stress-strain methods have made it possible to compute elastic constants for high-throughput databases of ordered structures, as well as individually tailored alloy compositions. However, this approach depends on predetermined or iteratively-chosen strain tensors. This poses a significant challenge for systematic, composition-dependent studies of disordered materials requiring structural models with randomized atomic positions and possessing low symmetry. Density functional perturbation theory (DFPT) based on density functional theory provides a compelling alternative to stress-strain methods: it allows for an unbiased determination of the elastic tensor, as well as access to local field data derived from the underlying general response function framework. Despite its intrinsic flexibility and efficiency, DFPT has seen limited application to the study of disordered systems. At the same time, there is a growing need for expanded quantum mechanical data to improve predictive modeling of complex disordered material properties. Here we present results for the rigid-ion and relaxed-ion elastic tensors computed using DFPT, for a comprehensive set of disordered structural refractory body-centered cubic (BCC) binary alloys of molybdenum (Mo), niobium (Nb), tantalum (Ta), and tungsten (W). For the first time, we have mapped the quantum-driven heterogeneity in elastic constants and derived mechanical properties—bulk modulus ($B$), shear modulus ($G$), Young’s modulus ($E$), and Poisson’s ratio ($\nu$)—and associated nuclear-relaxation fields at each disordered structure lattice site. The derived properties, together with Pugh’s ratio ($B/G$), Cauchy pressure and elastic anisotropy, are reported as a function of composition for all refractory binaries. The DFPT-computed elastic constant results for the refractory binary alloys at systematically-varied Mo, Nb, Ta, and W compositions are in excellent agreement with available experimental data.
@@ 1 Introduction
The elastic tensor [dieter1976mechanical] and its elastic constant components give a complete description, in terms of deformation, of material response under external stress within the elastic limit. The elastic tensor can therefore serve as a valuable quantitative probe of elastic strain effects associated with changes in chemical composition, and their impact on mechanical [liang2024first, marker2018effects, muzyk2011phase, huhn2014first, hill1952elastic], electronic [bouarissa2001mechanical, de2007ab], thermodynamic [laberge1995elastic, zhang2010effects], and structural [born1940stability, born1954dynamical, persson1999dynamical] properties in these materials.
@@
Refractory alloys show great potential to meet the demands for high-temperature materials due to their high melting points and ability to maintain mechanical properties at temperatures higher than current Ni-based superalloys [senkov2018development]. The refractory elements Nb, Ta, Mo, and W of groups V and VI comprise the BCC structural refractory metals family, characterized by high mechanical strength and melting temperatures [nieh1993recent]. Their structural refractory binary alloys MoNb, MoTa, MoW, NbTa, WNb and WTa form solid solutions of BCC crystal structure across their entire compositional space [bijjala2024understanding], and form an attractive starting point for designing more complex disordered alloys. However, elastic constant measurements are typically reported only for a limited number of alloy classes and compositions, and experiments often exhibit substantial disagreement even between values measured using the same technique and on the same material [davies1977transducer, beg1976study]. Multiple factors can influence differences in reported experimental values, such as impurities and measurement temperature [shang2007first, hu2021screening, borges2024local, giles2022machine]. Previous efforts have focused on the computational determination of elastic constants using density functional theory (DFT) for selected refractory binary alloys, using stress-strain methods [jiang2018first, mei2023predicting, ge2017elastic, bhandari2022predicting, wang2020first]. While these investigations have provided valuable insights, they were limited to a few binary systems or computed for a narrow compositional range. The computational methodologies employed in these studies were primarily validated using experimental data for pure elements. Since the electronic interactions in alloy systems are inherently more complex than in their elemental counterparts, the direct application of methods validated for pure elements to alloy systems may introduce significant uncertainties, potentially compromising the accuracy and practical relevance of the predicted elastic properties for these alloys. These challenges are compounded when modeling disordered materials [horton2025accelerated]. Here, we establish and validate a first-principles modeling framework based on DFT [hohenberg1964, kohn1965] for performing computationally-efficient studies of elastic tensor and derived properties of disordered alloys, validated for systematically-varied Mo, Nb, Ta, and W binary alloy compositions. DFT, as implemented in state-of-the-art plane wave pseudopotential periodic electronic structure codes such as Abinit [gonze2020abinit, romero2020abinit], provides direct access to the computation of energies and stresses for obtaining the elastic tensors of the materials, with only atomic structure and nuclear charges as inputs.
@@
where $\bm{\sigma}$ is the $6\times 1$ stress tensor, the $C_{ij}$ are the elastic constants, and $\bm{s}$ is the $6\times 1$ strain tensor in Voigt notation [simmons1971single].
In the traditional stress-strain approach, the elastic constants $C_{ij}$ are obtained by computing either the total energy or the stress of a crystal at a series of preselected strains [shang2007first, dal2016elastic, luo2012ab, shao2012temperature, steinle1999first]. The selection of strain sets can significantly influence the computed elastic constants, with variations exceeding 15% [de2015charting] in some cases. To mitigate this problem, high-throughput databases of ordered materials, such as the Materials Project [Jain2013], validate computed elastic constants by comparing results obtained from multiple strain sets [de2015charting]. In disordered alloys, such as the refractory binary alloys that are the focus of the present work, this approach is not practical due to the lack of crystal symmetry caused by intrinsic lattice strain.
An alternative and more general approach that avoids the need for tailored strains is based on analytical differentiation of the energy using linear response theory [baroni1987elastic, hamann2005metric, hamann2005generalized, hamann2005systematic]. In this method, the elastic tensor is obtained by expanding the total energy with respect to infinitesimal perturbation of the system using density functional perturbation theory (DFPT) [baroni2001phonons].
@@
DFPT for the elastic tensor was initially implemented for norm-conserving pseudopotentials (NCPPs) [Gonze1997a] and later extended to the projector augmented-wave (PAW) method by Martin et al. [martin2019projector]. DFPT has been successfully demonstrated using both NCPPs [soumelidou2017strain, stenuit2007first, bouamama2009ab] and PAW pseudopotentials [daoust2022longitudinal, daoust2021impact, ma2022investigation, zha2023alloying, peng2019tunable, jia2021origin]. Previous applications of the DFPT methodology have primarily focused on the study of piezoelectric materials [daoust2022longitudinal, daoust2021impact, ma2022investigation, zha2023alloying, peng2019tunable, jia2021origin, ko2022first, zhang2025deep, zha2023enhanced, wang2022band]. Despite its intrinsic advantages, DFPT for elastic properties has been applied to only a limited number of structural alloys, e.g. the ternary alloys $({\rm Al}_{x}{\rm Mg}_{1-x}){\rm Sc}$ [saha2014elastic] and alloys of aluminum nitride and rare-earth nitride [daoust2022longitudinal]. The present work represents the first demonstration of DFPT for computing the elastic properties of the structural refractory disordered binary alloys of Mo, Nb, Ta and W.
Within the DFPT framework, the elastic tensor can be evaluated as the rigid-ion elastic tensor, $C_{ij}^{rigid}$, or as the relaxed-ion elastic tensor $C_{ij}^{relaxed}$ [baroni1987elastic, hamann2005metric, hamann2005generalized, hamann2005systematic]. $C_{ij}^{rigid}$ is computed with the relaxed initial structure, and strain as the perturbation; $C_{ij}^{relaxed}$ is a correction to the rigid-ion elastic tensor, with both strain and atomic displacements (nuclear relaxations) as the perturbations. Both $C_{ij}^{rigid}$ and $C_{ij}^{relaxed}$ provide essential information about the elastic response of a material. The relaxed-ion elastic tensor is the quantity that will correspond most closely to experiment.
Comparison of the rigid-ion and relaxed-ion results allows an assessment of the accuracy of the rigid-ion approximation, as well as enabling the mapping of the nuclear relaxation field of a disordered alloy. We define the nuclear relaxation elastic tensor $C_{ij}^{\text{nuclear}}$ as the difference between the corresponding rigid-ion and relaxed-ion quantities [hamann2005systematic]:
@@
Non-zero nuclear-relaxation contributions in binary alloys are due to lattice distortions induced by the alloying elements, leading to forces between atoms in a finite strain state. Nuclear-relaxation contributions have previously been used to understand the effect of stress on piezoelectric coefficients and elastic constants by decomposing these properties into a purely electronic and nuclear-relaxation contributions resulting from the atomic rearrangements in the strained state [daoust2022longitudinal, daoust2021impact, ma2022investigation, ko2022first, zhang2025deep, zha2023alloying, zha2023enhanced, wang2022band]. In the study by Zhang et al. [zhang2006first] and in recent works by Zha et al. [zha2023alloying, zha2023enhanced], the nuclear-relaxation term of the piezoelectric coefficient was further decomposed into lattice point-specific relaxation contributions in order to elucidate the microscopic origin of the piezoelectric effect. In a recent study by Zhang et al. [zhang2025deep] specific atomic contribution to piezoelectric response in $\rm{(K,Na)NbO_{3}}$ was studied. Despite the methodological similarities between the computation of relaxed-ion elastic constants and relaxed-ion piezoelectric properties, the atom specific contributions from nuclear relaxation to the elastic constants have not been previously reported. In this study, for the first time, microscopic insights into the influence of alloying on the mechanical behavior of refractory binary alloys are provided through the decomposition of the nuclear relaxation contributions. This approach enables the investigation of the spatial variation of elastic constants and derived elastic moduli within the crystal structure of disordered refractory binary alloys. This allows a deeper understanding of how local atomic environments affect the overall mechanical response of these complex systems.
@@
In the present work, we utilize DFT+DFPT to compute the single-crystal, relaxed-structure elastic tensors of the six BCC refractory binary alloys of (Mo, Nb, Ta, W)—MoNb, MoTa, MoW, NbTa, MoTa, and MoW—as a function of composition. The disordered alloys are modeled as supercells constructed using the special quasi-random structures (SQS) method [zunger1990special]. The relaxed disordered structures are then used to compute the rigid-ion and relaxed-ion elastic tensors based on DFPT. We assess the effectiveness of DFPT in providing accurate elastic tensors through detailed comparison of elastic constants with available experimental data. The composition-dependent calculations serve as an important test of the ability of the SQS method to accurately model disorder in these alloys. Mechanical properties $B$, $G$, and $E$, the elastic anisotropy and $B/G$ and $C^{\prime}$ derived from the computed elastic moduli are reported as a function of composition for all refractory binaries. Computed relaxation fields for the Voigt-averaged elastic constants, shear modulus, bulk modulus, Young’s modulus, and Poisson’s ratio are presented for all binaries and discussed in terms of constituent element chemistry. The in-depth understanding of trends in elastic properties and anisotropy in the disordered refractory binary alloys gained in this work through the decomposition of elastic constants into electronic and nuclear contributions is expected to contribute to new microscopic alloy computational strategies for designing refractory alloys for high-temperature applications.
@@ 2 Computational Methods
We used the open-source, periodic plane wave pseudopotential electronic structure code Abinit [gonze2020abinit, romero2020abinit] for all DFT calculations, employing the Perdew, Burke, and Ernzerhof (PBE) generalized gradient approximation (GGA) exchange-correlation functional [perdew1996generalized], with corresponding PBE norm-conserving pseudopotentials [van2018pseudodojo]. We note that while meta-GGA methods such as r2SCAN [furness2020accurate] and parameterized DFT+$U$ [anisimov1991band] approaches have seen growing popularity for computing the properties of ordered systems with 1–3 elements [kingsbury2022performance], the corresponding pseudopotentials are not currently available, and the accuracy of these functionals for complex disordered systems with 5+ elements has yet to be assessed. Since the present work is intended as a starting point for scaling up to larger refractory alloy systems with increased chemical complexity, the PBE functional is utilized throughout this work, to ensure a uniform level of theory across all alloys.
@@
The plane waves for each system were expanded up to a kinetic energy of 30 Ha, with each structure converged until the total residual differences in the potential were $10^{-12}$ Ha, corresponding to convergence to $10^{-6}$ bohr for lattice constants, and atomic forces converged to $|\pm 0.01|$ mHa/bohr. Following careful convergence studies, we adopted a $\Gamma$-centered Monkhorst–Pack [monkhorst1976special] special $k$-point grid with a density of 50 $k$-points per Å for all systems.
@@
Initial benchmark studies were performed on bulk Mo, Nb, Ta, and W in order to determine appropriate computational parameters for the binary alloy calculations. We computed the lattice constants, bulk moduli, cohesive energies, elastic constants, and (100), (110), (111) surface energies for elemental Mo, W, Nb, and Ta, and compared our results with available experimental data, using three different pseudopotentials: the FHI norm-conserving pseudopotentials ($d$ valence electrons only) tabulated at the Abinit website [Abinit-website], pseudo-dojo [van2018pseudodojo] NCPPs ($s$, $p$ and $d$ valence electrons) available at [pseudodojo-website], and projector augmented-wave (PAW) pseudopotentials [blochl1994projector] ($s$, $p$ and $d$ valence electrons), tabulated at the Abinit website [Abinit-website]. Computed properties using the FHI NCPPs showed an average deviation of $\approx 1\%$ compared to available experiments, across all computed properties for Mo, Nb and Ta, but an average deviation of $\approx 9\%$ for W. For the pseudo-dojo NCPPs and PAW-based pseudopotentials, computed properties agreed well with experiment, with an average deviation of $\approx 1.5\%$ for all elements and across all properties. Detailed results are reported in [bijjala2025composition]. Given the comparable performance of the pseudo-dojo NCPP and PAW pseudopotentials, the pseudo-dojo NCPPs were used throughout this work. Working within the pseudo-dojo framework [van2018pseudodojo] also provides a foundation for the custom construction of pseudopotentials based on emerging functionals, for computationally-consistent modeling of alloys of growing complexity.